you are plannibg to make a series of deposits for an expense in the future. you will deposit 5,000 next year, 4,000 in five years abd 3,000 in eight years from today. interest rate is 4.50%, how much will you have in 10 years?
Answer #1

Ans $ 15691.29

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
a) FV = PV (1 + r )^n
FV = 5000*(1+4.5%)^9
FV = 7430.48
b) FV = PV (1 + r )^n
FV = 4000*(1+4.5%)^5
FV = 4984.73
c) FV = PV (1 + r )^n
FV = 3000*(1+4.5%)^2
FV = 3276.08
TOTAL FV = 7430.48 + 4984.73 + 3276.08
15691.29